OpenStudy (anonymous):

Is there an easy way to find the C and k of the Big O notation in Discrete Mathematics?

OpenStudy (anonymous):

not sure what C and k are as i recall \[f(x)=O(g(x))\] if \[\lim_{x\to \infty}\frac{f(x)}{g(x)}=M\]

OpenStudy (anonymous):

hmmm interesting. I've never seen that equation. But here's what I know, "we say that f(x) is O(g(x)) if there are constants C and k such that \[|f(x)| \le C|g(x)|\] whenever x >k

OpenStudy (anonymous):

I understand the concept but I'm having a hard time figuring out how to find the C and k in a simple way

OpenStudy (anonymous):

ok they say the same thing really divide by \(|g|\) and they are equivalent

OpenStudy (anonymous):

i seriously doubt there is one method to solve for all \(f\) and \(g\)

OpenStudy (anonymous):

Oh ok. thanks! :)

OpenStudy (anonymous):

Another question! "Show that \[7x^2\] is \[Ox^3\]"

OpenStudy (anonymous):

in the solution, when x >7, they had 7x^2 < x^3

OpenStudy (anonymous):

Is it a trial and error to come up with the value of k?

OpenStudy (anonymous):

i must have confused big o notation with little o notation i would not have said that was correct, so i was wrong in any case i guess big O means the function is bounded above by the other function, not that they grew at the same rate in any case you can do it by trial and error it is clear that \(x^3\) grows faster than \(x^2\) right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

or rather grows faster than \(7x^2\)

OpenStudy (anonymous):

well then there is not much to this if \(x=7\) then \(7x^2=7^3\) and \(x^3=7^3\) they are equal when \(x=7\) if \(x>7\) then \(7x^2<x^3\)

OpenStudy (anonymous):

we can try another if you like, but this example seems pretty straight forward

OpenStudy (anonymous):

Oh whew I didn't realize that it was that easy lol. Sure! we could try another one. I found 2 problems. 1) Show that x^2 + 2x +1 is O(x^2) 2) Show that 4x^2 +2x +! is 0(x^4)

OpenStudy (anonymous):

*2) 4x^2 +2x +1

OpenStudy (anonymous):

for what value of \(x\) is \(x^2+2x+1>x^2\) ?

OpenStudy (anonymous):

actually \(x^2+2x+1=o(x^2)\) as the limit is one

OpenStudy (anonymous):

oh maybe i got this one backwards no matter \[x^2+2x+1\leq C(x^2)\]for some value of \(C\) we could pick a C less than 1, say \(\frac{1}{2}\)

OpenStudy (anonymous):

then since \[\lim_{x\to \infty}\frac{x^2+2x+1}{x^2}=1\] oh crap C>1 in any cases, find some number for which \[\frac{x^2+2x+1}{x^2}<2\]

OpenStudy (anonymous):

probably a small number will do, like 3